Analyze Panera Bread Company Panera Bread Company (PNRA) operates over 2,000 bakery-cafe locations throughout the United...

  1. Analyze Panera Bread Company

    Panera Bread Company (PNRA) operates over 2,000 bakery-cafe locations throughout the United States and Canada and serves over 9 million customers per week. Panera’s operations are divided into the following segments:

    • Company-Operated Bakery-Cafes
    • Franchised Bakery-Cafes
    • Fresh Dough and Other Products

    The Fresh Dough and Other Products segment supplies fresh dough, produce, tuna, and other products to the company-operated and franchised cafes. Recent data (in millions) for each of these segments are as follows:

    Company-Operated Cafes Franchised Cafes Fresh Dough
    Revenues $2,434 $155 $408
    Operating income 398 150 28
    Invested assets 943 18 87

    a. Determine the profit margin for each segment. Round to one decimal place.

    Profit margin
    Company-Operated %
    Franchised %
    Fresh Dough %

    b. Determine the investment turnover for each segment. Round to two decimal places.

    Investment
    Turnover
    Company-Operated
    Franchised
    Fresh Dough

    c. Use the DuPont formula to determine the return on investment for each segment. Round to one decimal place.

    Return on
    Investment
    Company-Operated %
    Franchised %
    Fresh Dough %
Answer #1

a. Determine the profit margin for each segment. Round to one decimal place.

Profit margin
Company-Operated 398/2434 = 16.4%
Franchised 150/155 = 96.8%
Fresh Dough 28/408 = 6.9%

b. Determine the investment turnover for each segment. Round to two decimal places.

Investment
Turnover
Company-Operated 2434/943 = 2.58
Franchised 155/18 = 8.61
Fresh Dough 408/87 = 4.69

c. Use the DuPont formula to determine the return on investment for each segment. Round to one decimal place.

Return on
Investment
Company-Operated 16.4*2.58 = 42.3%
Franchised 96.8*8.61 = 833.4%
Fresh Dough 6.9*4.69 = 32.4%